OpenStudy (anonymous):

Not sure how they got to this answer in my sol. manual. The integral of 2/(v^2-v) dv = 2ln[(v-1)/v]

OpenStudy (anonymous):

\[\int\limits\limits_{}^{} 2/(v^2-v) dv = 2\ln ((v-1)/v)\]

OpenStudy (anonymous):

take v common in the denominator, ull get 2/v(v-1) dv. Now you can take partial fractions. Understood?

OpenStudy (dumbcow):

partial fractions v^2 -v = v(v-1) A/v + B/v-1 = A(v-1) + Bv / v(v-1) = 2/v(v-1) (A+B)v - A = 2 -A = 2 --> A = -2 A+B = 0 --> B=2 integral -2/v + 2/v-1 -2ln(v) +2ln(v-1) +C = 2ln(v-1/v) +C

OpenStudy (anonymous):

you need to present v^2-v=v(v-1) after that: A/v + B/(v-1)=... find A & B now your integral could be presented as: \[=\int\limits_{ }^{}-2dv/v + \int\limits_{ }^{}2dv/(v-1)=...\] i think you can take it from here...?

OpenStudy (anonymous):

Yes! Thank you!